OpenStudy (anonymous):

A small publishing company is planning to publish a new book. The production costs will include one-time fixed costs (such as editing) and variable costs (such as printing). There are two production methods it could use. With one method, the one-time fixed costs will total $15,192 , and the variable costs will be $21.50 per book. With the other method, the one-time fixed costs will total $45,690, and the variable costs will be $11.75 per book. For how many books produced will the costs from the two methods be the same?

OpenStudy (anonymous):

cost = c books made =b method 1: c=21.5b+15192 method 2: c=11.75b+45690 to find where they are equal 21.5b+15192=11.75b+45690 b=3128 books